# MS-Glance: Non-semantic context vectors and the applications in supervising image reconstruction

### v1

@@ 1 Introduction
In recent years, research in computer vision (CV) has increasingly incorporated global semantic information, inspired by the human visual perception process [23] where understanding is driven by high-level semantic cues rather than focusing on individual pixels. For example, the Vision Transformer (ViT) [11], a prominent neural network architecture, captures long-range dependencies within an image and generates powerful feature representations through vision-language pretraining[60]. In the field of image reconstruction, advanced models have also achieved success in incorporating semantic information. For instance, in super-resolution, architectures such as U-Net utilize multiple layers of convolutional and downsampling operations to progressively extract high-level semantic information [61, 25, 80]. Besides, in undersampled MRI reconstruction, some researchers utilize semantic segmentation networks to assist in the reconstruction process [28, 14, 55].
However, in human vision, semantic conception[67] is not the only way to portray the recognition; it can also be interpreted as a reconstruction procedure of the scene that doesn’t involve semantic understanding[48, 54, 63, 49]. A pioneering study[54] shows that human vision gleans a large amount of meaningful information from a single glance. Several experimental studies followed[63, 49], suggesting that recognition of the real world may start from encoding global configuration since no perception of individual objects or detailed features can be made in such a short time. A representative computational model, Spatial Envolope[48], models the structure of real-world scenes by a set of perceptual properties including naturalness, openness, roughness, ruggedness, and expansion. Further, Michelle Greene et al. [19] experimentally shows that statistical and structural cues extracted from very brief (e.g. 19 ms and masked) exposures allow for above-chance categorization of scenes, verifying the Spatial Envelope model and the existence of a reconstruction-based recognition.
@@ 2.1 Image non-semantic information
While semantic information has been extensively utilized in various computer vision tasks, research on non-semantic information of images, such as structural and layout features, remains limited. Cao et al. [1] introduced the concept of Non-Semantic Facial Parts (NSFP), which identifies the most discriminative patches for face recognition and retrieval. However, their method relies on predefined features like SIFT, limiting the performance. Murrugarra et al. [46] proposed non-semantic transfer from attributes that may belong to different domains, but focuses solely on texture, which is only a small subset of non-semantic information. In autonomous driving, Anas et al. [3] represented point clouds with non-semantic features for environment interpretation, localization, and mapping; yet their work is specific to point clouds rather than images. In NeRF, Xie et al. [75] applied SSIM on reorganized random image patches. Still, their work is also application-specific and relies on the original pixel-based SSIM computation with a local focus center.
@@ 2.2 Image reconstruction
Image reconstruction [78, 29, 69] in CV refers to the task of recovering or restoring an image from incomplete, corrupted, or undersampled data. Here, we primarily introduce two image reconstruction tasks: Implicit Neural Representation (INR) and undersampled MRI reconstruction.
@@ 2.2.1 Implicit Neural Representation (INR)
INR parameterizes a given discrete image $f$ in a continuous fashion. The image defines a dataset
$\mathcal{D}=\{(\mathbf{x}_{i},f(\mathbf{x}_{i}))_{i}\},f:\mathbb{R}^{2}\mapsto\mathbb{R}^{3}$
of pixel coordinates $\mathbf{x}_{i}=(x_{i},y_{i})$ associated with their RGB colors $f(\mathbf{x}_{i})$.
There has been various research on INR for better image fitting, for example, Sinusoidal Representation Networks (SIREN) [64] utilizes periodic activation functions for INR to capture intricate natural signals and their derivatives. INR has also been used for image compression [68, 12], segmentation [35], and super-resoultion [4, 47]. In this work, we focus on its basic image-fitting performance for a proof-of-concept. One major application of implicit neural representation is a series of works [45, 71, 81] on 3D reconstruction. Instead of using explicit representation(e.g., voxels), NeRf [45] encodes the representation of 3D object/scene into MLPs. Taking the 3D location and camera view angle as input, MLP can output the density and color to finally accomplish the 3D reconstruction task.
@@ 2.3 Loss functions for comparing images
In tasks related to image comparing, various network architectures have been extensively explored however research on loss functions remains relatively underdeveloped. There are primarily two categories of loss functions commonly used: (1) pixel-wise loss functions such as L1 [31] and L2 loss [76]. These types of loss functions operate at the level of individual pixels or elements in the output and target images. (2) global loss functions such as Gram loss [16], adversarial loss (GAN) [18], perceptual loss[82], and SSIM loss [73, 75]. In contrast to pixel-wise loss functions, global loss functions consider the relationships and structures across the entire image. There has been research [84] that indicates the quality of the results improves significantly with better loss functions, empirically a combination of pixel-wise loss and global loss. However, there is currently no effective loss function that can capture information about non-semantic image context.
@@ 4.1 Image Fitting with INR
We further evaluate the performance of MS-Glance compared to other methods during the SIREN regression process. Using a classic RGB image of shape $512\times 512$, Astronaut [^6] https://scikit-image.org/docs/stable/api/skimage.data.html, we plot PSNR and SSIM at each step of the SIREN regression, as shown in Fig. 3. MS-Glance consistently outperforms existing loss functions throughout the entire training process. Additional qualitative results using Astronaut are provided in the supplementary materials.
@@ 4.2 Undersampled MRI reconstruction
Model Implementation. DRDN’s input is a 2-channel undersampled MRI image and the output is a 2-channel reconstructed image. It stacks densely connected atrous layers and thus has a large receptive field while preserving image details. We use the official implementation [^7] https://github.com/bbbbbbzhou/DuDoRNet of DRDN and change its recurrent number into 3 due to the GPU memory limit. DRDN is trained with an Adam optimizer and a learning rate of 0.0001 for 40 epochs for IXI and 20 epochs for FastMRI, and the model checkpoints used for evaluation are the one that produces the best PSNR on the validation set.
@@
Loss implementation. All compared losses are scaled with a coefficient of 0.1 to match the order of magnitudes of the L2 losses before combining with L2. For our MS-Glance and Global Glance, we use air prior when selecting $\mathcal{V}_{Global}$ and set $\sigma$ as 0.01.
Result. Fig. 5. presents the qualitative results. The red box highlights a portion of the brain cortex, indicated by the red arrow. Among all, only MS-Glance and LPIPS successfully reconstruct the structure from the blurry input. However, LPIPS tends to introduce hallucinations, such as the vertical line on the right of the ‘+LPIPS’ text which is not in the Ground Truth. The quantitative results are provided in Table 2 and show similar behaviors to INR image fitting. MS-Glance achieves the best PSNR among all.
@@ 5 Ablation studies
5.1 The selection of $n\cdot m$ and $n_{g}\cdot m_{g}$
5.2 Effectiveness of the rule-based pixel selection
The construction of $\mathcal{V}_{Global}$ can integrate perceptual prior, and we have shown an example of using MRI air prior. To validate the effectiveness of this prior, we evaluate the performance of training undersampled MRI networks with and without air prior. The experiments are carried out on the FastMRI dataset under two acceleration rates. Table 5 presents the performance of MS-Glance with and without the MRI air prior. This validates its beneficial effect.
@@ 6.1 MS-Glance and SSIM
SSIM [73] computes image similarity in a window-based way and emphasizes the center pixel with a Gaussian kernel. SSIM is the product of three statistics: luminance $l$, contrast $c$, and structure $s$:
@@
where $\mu_{x},\mu_{y}$ represents the pixel sample mean of two windows $x$ and $y$, $\sigma_{x},\sigma_{y}$ represents the variance of $x,y$ and $\sigma_{xy}$ represents covariance of $x,y$. The SSIM structure term resembles the formulation of the Glance Index. However, the construction of Global Glance vectors gathers pixels across the image, rather than in local windows like SSIM. Additionally, when we combine the $l$ and $c$ of SSIM with our Glance Index, we witness a performance drop. We argue that the non-semantic image context we try to capture is highly correlated with image structure and thus is better measured by our proposed Glance Index unitarily. Moreover, MS-Glance does not assume locality or center for each window, since the image context is statistics-based and does not have a focus center. Ablation studies about the kernel selection can be found in the supplementary.
@@ 6.2 Future Work
MS-Glance incorporates multi-scale, non-semantic image context into supervised image reconstruction. It can be deployed in other low-level vision tasks such as super-resolution, deblurring, and compression. Also, the stochasticity property of Global Glance allows for integrating more perception priors. Future work includes combining MS-Glance with other explicit appearance priors[41, 40] and learned implicit priors such as pixel correlation[30].
@@ 7 Conclusion
In this paper, we show how to leverage the non-semantic image context, which can be captured by human vision, for supervising image reconstruction networks. We propose MS-Glance, a novel biologically inspired multi-scale descriptor of non-semantic image context and its loss form. We demonstrate its effectiveness in improving image reconstruction through two tasks: INR fitting and undersampled MRI reconstruction. Finally, extensive experiments on a wide range of natural and medical datasets show that the MS-Glance loss outperforms existing loss functions used in image restoration.
8 Supplementary Materials
Here, we show some additional results mentioned in the main paper: qualitative results on fitting the Astronaut image and ablation studies on the Glance’s window kernel and distance measure. We also add additional details on MS-Glance’s implementation and the network architecture of DRDN, which we use for undersampled MRI reconstruction experiments.
8.1 More qualitative results of Astronaut
Astronaut is a color image of the astronaut Eileen Collins. In Figure 6, we compare the step-wise reconstruction of Astronaut by SIREN and SIREN+MSGlance. The reconstructed images and the corresponding SSIM error maps are visualized. MS-Glance reconstructs the image details faster (the blue boxes in step 40) and ends up with a finer reconstruction (the blue boxes in step 500).
Figure 6: Step-wise reconstruction of the example image, Astronaut
8.2 More ablation studies
8.2.1 Uniform Kernel and Gaussian Kernel
The uniform window kernel is a key distinction between the Glance Index Measure and methods like SSIM and S3IM. To conduct a comprehensive ablation study, we replaced our uniform kernel with their Gaussian kernel on both tasks. For MRI reconstruction, we use the IXI dataset under two acceleration rates. For INR fitting, we use the Coco dataset.
Table 6 validates that compared with the Gaussian kernel, our uniform kernel not only stabilizes training but also enhances performance. Initially, when we applied the Gaussian kernel to MS-Glance, it caused the loss function to produce NaN values. To mitigate this, we temporarily switched to a standard $L_{p}$ loss during NaN iterations. However, a significant number of steps still resulted in NaN values. To further investigate, we decomposed MS-Glance. We observed that both Local Glance with Gaussian and MS-Glance with Gaussian led to approximately 60% NaN loss, contributing to the large performance degradation. While the Global Glance with Gaussian’s training remained stable, it also experienced a performance decline.
Undersampled MRI reconstruction INR fitting
5x 7x
PSNR (dB) SSIM PSNR (dB) SSIM PSNR (dB) SSIM
MS-Glance 31.434 0.9535 30.865 0.9485 35.249 0.9493
MS-Glance - Gaussian 29.497 0.9323 28.699 0.9229 35.099 0.9469
Global Glance 31.122 0.9524 30.711 0.9485 34.843 0.9439
Global Glance - Gaussian 31.104 0.9524 30.622 0.9483 34.827 0.9441
Local Glance 31.346 0.9537 30.813 0.9483 35.004 0.9463
Local Glance - Gaussian 29.487 0.9315 28.573 0.9208 34.830 0.9430
Table 6: Ablation study of the window kernel.
8.2.2 Glance Index Measure and SSIM
While we compare our method with SSIM loss in all experiments, we also highlight the connection between the Glance Index Measure and SSIM, which is discussed in detail in the main paper. In this section, we provide additional experimental results to compare the performance of the Glance Index Measure against SSIM. As mentioned in the main paper, the structural term of SSIM computes covariance similarly to how the Glance Index Measure operates. However, SSIM also incorporates luminance ($l$) and contrast ($c$) terms. To account for this, we extend our Glance Index Measure by integrating the computation of $l$ and $c$, multiplying them with the original Glance Index Measure. We tested this modified approach across both tasks.
We perform the evaluation on both tasks. For MRI reconstruction, we use the IXI dataset under two acceleration rates. For INR fitting, we use the Coco dataset. Table 7 demonstrates the effectiveness of the Glance Index Measure, particularly in global scenarios. The current Glance Index Measure shows that MS-Glance and Global Glance remain superior. However, the Local Glance enhanced with $l$ and $c$ exhibits improved performance, especially in SSIM computations. This improvement is expected, as it directly optimizes a term similar to SSIM itself. Additionally, we explored combining the original Global Glance design with the new Local Glance incorporating $l$ and $c$, with results shown in the last row. This approach, however, did not perform as well as the original MS-Glance design, suggesting a conflict between the two approaches.
Undersampled MRI reconstruction INR fitting
5x 7x
PSNR (dB) SSIM PSNR (dB) SSIM PSNR (dB) SSIM
MS-Glance 31.434 0.9535 30.865 0.9485 35.249 0.9493
+ $l$ and $c$ of SSIM 31.035 0.9523 30.660 0.9483 35.131 0.9484
Global Glance (a) 31.122 0.9524 30.711 0.9485 34.843 0.9439
+ $l$ and $c$ of SSIM 31.055 0.9523 30.653 0.9479 34.714 0.9426
Local Glance 31.346 0.9537 30.813 0.9483 35.004 0.9463
+ $l$ and $c$ of SSIM (b) 31.425 0.9557 30.939 0.9519 34.913 0.9468
Combination of (a) and (b) 30.953 0.9516 30.701 0.9492 35.242 0.9496
Table 7: The effect of Glance Index Measure and SSIM to MS-Glance.
8.3 Additional Details
8.3.1 Implementation of MS-Glance
In the Global Glance process, we randomly select pixels and shuffle them 10 times, resulting in more Glance vectors for computing the Global Glance Index Measure. As shown in Table 8, shuffling leads to a slight improvement in performance. The experiments are carried out on the Coco dataset.
Shuffle times 1 5 10
PSNR 35.202 35.257 35.249
SSIM 0.9489 0.9491 0.9493
Table 8: The effect of the shuffle time on INR fitting.
8.3.2 Architecture of DRDN
We choose DRDN as the network for undersampled MRI reconstruction. Its strong performance has been validated by their original experiments and many recently established works[7, 65]. DRDN[85] customizes the local and global structure design for the MRI reconstruction task. It uses a Squeeze-and-excitation Dilated Residual Dense Block (SDRDB) as the backbone. The main diagram is shown in Figure 7.
Figure 7: The diagram of DRDN[85].
Globally, DRDN consists of an initial feature extraction module (two sequential 3 × 3 convolution layers), multiple SDRDBs followed by global feature fusion (a concatenation operation for all SDRDBs’ output), and global residual learning enhanced by a Squeeze-and-excitation on the residual branches.
Figure 8: The component of each SDRDB.
The structure in each SDRDB is shown in Figure. In each SDRDB, there are four densely connected atrous convolution layers, local feature fusion, Squeeze-and-Excitation, and local residual learning.

### v2

# MS-Glance: Bio-Inspired Non-semantic Context Vectors and their Applications in Supervising Image Reconstruction

@@ 1 Introduction
Computer vision (CV) has increasingly incorporated global semantic information, inspired by the human visual perception process [25] where understanding is driven by high-level semantic cues rather than focusing on individual pixels. For example, the Vision Transformer (ViT) [11], a prominent neural network architecture, captures long-range dependencies within an image and generates powerful feature representations through vision-language pretraining[62]. In image reconstruction, advanced models have also successfully incorporated semantic information. For instance, in super-resolution, architectures such as U-Net utilize multiple layers of convolutional and downsampling operations to progressively extract high-level semantic information [63, 27, 82]. Besides, in undersampled MRI reconstruction, some researchers utilize semantic segmentation networks to assist in the reconstruction process [30, 14, 57]. However, in human vision, semantic conception[68] is not the only way to portray the recognition; it can also be interpreted as a reconstruction procedure of the scene that doesn’t involve semantic understanding[50, 56, 65, 51], demonstrated in Fig.1. A pioneering study[56] shows that human vision gleans a large amount of meaningful information from one single glance. Experimental studies followed[65, 51], suggesting that recognition of the real world may start from encoding global configuration since no perception of individual objects or detailed features can be made in such a short time. A representative computational model, Spatial Envolope[50], models the structure of real-world scenes by a set of perceptual properties including naturalness, openness, roughness, ruggedness, and expansion. Further, Michelle Greene et al. [21] experimentally shows that statistical and structural cues extracted from very brief (e.g. 19 ms and masked) exposures allow for above-chance categorization of scenes, verifying the Spatial Envelope model and the existence of a reconstruction-based recognition.
@@ 2.1 Image non-semantic information
While semantic information has been extensively utilized in various CV tasks, research on non-semantic information of images, such as structural and layout features, remains limited. Cao et al. [1] introduced the concept of Non-Semantic Facial Parts (NSFP), which identifies the most discriminative patches for face recognition and retrieval. However, their method relies on predefined features like SIFT, limiting the performance. Murrugarra et al. [48] proposed non-semantic transfer from attributes that may belong to different domains, but focuses solely on texture, which is only a small subset of non-semantic information. In autonomous driving, Anas et al. [3] represented point clouds with non-semantic features for environment interpretation, localization, and mapping; yet their work is specific to point clouds rather than images. In NeRF, Xie et al. [76] applied SSIM on reorganized random image patches. Still, their work is also application-specific and relies on the original pixel-based SSIM computation with a local focus center.
@@ 2.2 Image reconstruction
Image reconstruction [80, 31, 70] in CV refers to recovering or restoring an image from incomplete, corrupted, or undersampled data. Here, we introduce one reconstruction method and one reconstruction task: Implicit Neural Representation (INR) and undersampled MRI reconstruction.
@@ 2.2.1 Implicit Neural Representation (INR)
INR is a neural network-based continuous image representation. It parameterizes a field in a coordinate-based manner. Various research on INR explored better image fitting methods, for example, Sinusoidal Representation Networks (SIREN) [66] utilizes periodic activation functions for INR to capture intricate natural signals and their derivatives. INR has also been used for image compression [69, 12, 19], segmentation [37], and super-resoultion [4, 49]. Other major application of implicit neural representation is a series of works [47, 72, 83, 15] on 3D reconstruction. Instead of using explicit representation(e.g., voxels), NeRf [47] encodes the representation of 3D object/scene into MLPs. In this work, we focus on improving INR’s image-fitting performance, potentially applicable to various related tasks.
@@ 2.3 Loss functions for comparing images
In tasks related to image comparing, various network architectures have been extensively explored while loss functions remain relatively under-developed. There are primarily two categories of loss functions commonly used: (1) pixel-wise loss functions such as L1 [33] and L2 loss [77]. These types of loss functions operate at the level of individual pixels or elements in the output and target images. (2) global loss functions such as Gram loss [17], adversarial loss (GAN) [20], perceptual loss[84], and SSIM loss [74, 76]. In contrast to pixel-wise loss functions, global loss functions consider the relationships and structures across the entire image. There has been research [86] that indicates the quality of the results improves significantly with better loss functions, empirically a combination of pixel-wise loss and global loss. However, there is currently no effective loss function that can capture information about non-semantic image context.
@@ 4.1 Image Fitting with INR
We further visualize the SIREN regression process supervised by MS-Glance and other methods. Using a classic RGB image of shape $512\times 512$, Astronaut [^6] https://scikit-image.org/docs/stable/api/skimage.data.html, we plot PSNR and SSIM at each step of the SIREN regression, as shown in Fig. 3. MS-Glance consistently outperforms existing loss functions throughout the entire training process. Additional qualitative results using Astronaut are provided in the supplementary materials.
@@ 4.2 Undersampled MRI reconstruction
Model Implementation. DRDN’s input is a 2-channel undersampled MRI image and the output is a 2-channel reconstructed image. It stacks densely connected atrous layers and thus has a large receptive field while preserving image details. We use the official implementation [^7] https://github.com/bbbbbbzhou/DuDoRNet and change its recurrent number into 3 due to the GPU memory limit. DRDN is trained with an Adam optimizer and a learning rate of 0.0001 for 40 epochs for IXI and 20 epochs for FastMRI, and the model checkpoints used for evaluation are the one that produces the best PSNR on the validation set.
@@
Loss implementation. All compared losses are scaled with a coefficient of 0.1 to match the order of magnitudes of the L2 losses before combined with L2. For our MS-Glance and Global Glance, we use air prior when selecting $\mathcal{V}_{Global}$ and set $\sigma$ as 0.01.
Result. Fig. 5 presents the qualitative results. The red box highlights a portion of the brain cortex, indicated by the red arrow. Among all, only MS-Glance and LPIPS successfully reconstruct the structure from the blurry input. However, LPIPS tends to introduce hallucinations, such as the vertical line on the right of the ‘+LPIPS’ text which is not in the Ground Truth. The quantitative results are provided in Table 2 and show similar behaviors to INR image fitting. MS-Glance achieves the best PSNR among all.
@@ 5 Ablation studies
5.1 The selection of $n\cdot m$ and $n_{g}\cdot m_{g}$
5.2 Effectiveness of the rule-based pixel selection
The construction of $\mathcal{V}_{Global}$ can integrate perceptual prior, and we have shown an example of using MRI air prior. To validate the effectiveness of MRI air prior, we evaluate the performance of training undersampled MRI networks with and without it. The experiments are carried out on the FastMRI dataset under two acceleration rates. Table 5 presents the performance of MS-Glance with and without the MRI air prior. This validates its beneficial effect.
@@ 6.1 MS-Glance and SSIM
SSIM [74] computes image similarity window-wise and emphasizes the center pixel with a Gaussian kernel. Given two windows $x$ and $y$, their similarity is the product of three statistics: luminance $l$, contrast $c$, and structure $s$:
@@
where $\mu_{x},\mu_{y}$ are the mean of $x$ and $y$, $\sigma_{x},\sigma_{y}$ are the variance of $x,y$ and $\sigma_{xy}$ is the covariance of $x,y$. $s$ resembles the formulation of the Glance Index. However, the construction of Global Glance vectors gathers pixels across the image, rather than in local windows like SSIM. Additionally, when we combine the $l$ and $c$ of SSIM with our Glance Index, we witness a performance drop. We argue that the non-semantic image context we try to capture is highly correlated with image structure and thus is better measured by our proposed Glance Index unitarily. Moreover, MS-Glance does not assume locality or center for each window, since the image context is statistics-based and does not have a focus center. Ablation studies about the kernel selection can be found in the supplementary.
@@ 6.2 Future Work
MS-Glance incorporates multi-scale, non-semantic image context into supervised image reconstruction. It can be deployed in other low-level vision tasks such as super-resolution, deblurring, and compression. Also, the stochasticity property of Global Glance allows for integrating more perception priors. Future work includes combining MS-Glance with other explicit appearance priors[43, 42] and learned implicit priors such as pixel correlation[32, 78].
@@ 7 Conclusion
We have shown how to leverage the non-semantic image context, which can be captured by human vision, for supervising image reconstruction networks. We propose MS-Glance, a novel bio-inspired multi-scale descriptor of non-semantic image context and its loss form. We demonstrate its effectiveness in improving image reconstruction through two settings: INR fitting and undersampled MRI reconstruction. Finally, extensive experiments on both natural and medical datasets show that the MS-Glance loss outperforms existing loss functions used in image restoration.
8 Acknowledgement
This work is supported by Natural Science Foundation of China under Grant 62271465, Suzhou Basic Research Program under Grant SYG202338, and Open Fund Project of Guangdong Academy of Medical Sciences, China (No. YKY-KF202206).